OpenStudy (anonymous):

Evaluate the expression described below if the variable given is 9. Two times the difference of one-third of a number and three.

OpenStudy (anonymous):

2*((1/3)X-3) so start with X=9, (1/3)9 = 3: 2*(3-3) then 3-3 is 0, 2*0: finally you see that 0 times anything is just 0... So your answer would be 0

OpenStudy (anonymous):

Thank you! @sschroeder

OpenStudy (anonymous):

np